# Response Matrix Estimation in Unfolding Differential Cross Sections

### v1

@@ Abstract
The unfolding problem in particle physics is to make inferences about the true particle spectrum based on smeared observations from a detector. This is an ill-posed inverse problem, where small changes in the smeared distribution can lead to large fluctuations in the unfolded distribution. The forward operator is the response matrix which models the detector response. In practice, the forward operator is rarely known analytically and is instead estimated using Monte Carlo simulation. This raises the question of how to best estimate the response matrix and what impact this estimation has on the unfolded solutions. In most analyses at the LHC, response matrix estimation is done by binning the true and smeared events and counting the propagation of events between the bins. However, this approach can result in a noisy estimate of the response matrix, especially with a small Monte Carlo sample size. Unexpectedly, we also find that the noise in the estimated response matrix can inadvertently regularize the problem. As an alternative, we propose to use conditional density estimation to estimate the response kernel in the unbinned setting followed by binning this estimator. Using simulation studies, we investigate the performance of the two approaches.
@@
inverse problem , forward operator , conditional density estimation , Large Hadron Collider , Tikhonov regularization , D’Agostini iteration
@@ 1 Introduction
The unfolding problem arises in the statistical analysis of data at the Large Hadron Collider (LHC) at CERN. In LHC experiments, it is often of interest to learn the distribution (or differential cross section) $f$ of some physical quantity measured by the detectors, e.g., the energy, mass or production angle of the particles produced in the proton-proton collisions. However, due to the finite resolution of the detectors, we can only observe a smeared version of the truth, which results in a smeared measured distribution $g$ (Figure 1). Unfolding refers to the process of inferring $f$ from $g$ (for details, see, e.g., [1]).
@@
However, in practice, the response matrix $\bm{K}$ is usually not known analytically and needs to be estimated using detector simulations. This raises the question of how to best estimate the response matrix and what impact this estimation has on the unfolded solutions. The main contribution of this work is to introduce a new way to estimate the response matrix by first considering response kernel estimation in the unbinned setting. Additionally, we show that there are nontrivial, and sometimes unexpected, interactions between the response matrix estimators and the quality of the unfolded solutions.
@@ 2.1 Discretization
$p_{X}(x)=\frac{f(x)}{\lambda(T)},\;\;p_{Y}(y)=\frac{g(y)}{\mu(S)},$
@@ 2.2 Unfolded point estimation
$\bm{L}^{-1}\bm{y}$ $=\bm{L}^{-1}\bm{K}\bm{\lambda}+\bm{\eta}.$ (4)
Since $\bm{\eta}\sim N(\bm{0},\bm{L}^{-1}\bm{\Sigma}(\bm{L}^{-1})^{\top})$ and $\bm{L}^{-1}\bm{\Sigma}(\bm{L}^{-1})^{\top}=\bm{I}$, it follows that $\bm{\eta}\sim N(\bm{0},\bm{I})$. Therefore, since $\bm{\mu}$ is accurately estimable using $\bm{y}$ and the Cholesky decomposition always exists for the covariance matrix, we may replace $\bm{K}\hookrightarrow\bm{L}^{-1}\bm{K}$ and assume without loss of generality that $\bm{\epsilon}\sim N(\bm{0},\bm{I})$. Under this assumption and assuming that $\bm{K}$ has full column rank, the least squares estimator of $\bm{\lambda}$ is
$\hat{\bm{\lambda}}_{ls}=(\bm{K}^{\top}\bm{K})^{-1}\bm{K}^{\top}\bm{y}.$ (5)
@@
$\mathbb{E}[\hat{\bm{\lambda}}_{ls}]=(\bm{K}^{\top}\bm{K})^{-1}\bm{K}^{\top}\bm{K}\bm{\lambda}=\bm{\lambda}.$
@@
$\hat{\bm{\lambda}}_{tik}$ $=\text{argmin}_{\bm{\lambda}}\left\{\|\bm{y}-\bm{K}\bm{\lambda}\|_{2}^{2}+\delta\|\bm{\lambda}\|^{2}_{2}\right\}$ (6)
$=(\bm{K}^{\top}\bm{K}+\delta\bm{I})^{-1}\bm{K}^{\top}\bm{y},$
@@ 3.2.1 Kernel regression method
Estimating the conditional density can be reduced to a non-parametric regression problem. Let $K(\cdot)$ be a symmetric non-negative density function and let $h$ denote a bandwidth. Denote $f(y|x):=\mathbb{E}[K_{h}(Y-y)|X=x]$, where $K_{h}(\cdot)=K(\cdot/h)/h$. Then, by Taylor expansion,
$f(y_{0}|x)$ $=\int K_{h}(y^{\prime}-y_{0})p_{Y|X}(y^{\prime}|x)dy^{\prime}$
@@ 3.2.2 Local linear method
This yields the local linear estimate $\hat{p}_{ll}(y|x)=\hat{f}_{h_{1},h_{2}}(y|x)$. Note that if we set $b=0$ in the objective function (9), then the solution to the minimization problem recovers the kernel conditional density estimate $\hat{p}_{ks}(y|x)$. Compared to the kernel method, one attractive feature of the local linear method is its design adaptive property which states that asymptotically the bias of $\hat{p}_{ll}(y|x)$ does not depend on the marginal distribution of $X$ [10]. This can be favorable in many situations, for example, when the marginal distribution is falling steeply over many orders of magnitude. Also, more importantly, we do not know the distribution of $X$ in the first place.
@@ 3.2.3 Bandwidth selection
There are many existing methods for picking $h_{1}$ and $h_{2}$. One of the simplest and computationally most efficient ways is to find the optimal bandwidths by assuming the normality of both the conditional and marginal distributions [12, 13]. Although it relies on the normality assumption, this can work reasonably well even in cases where normality does not hold.
@@ 3.2.5 Location-scale model
$p_{Y|X}(y|x)=\frac{1}{\sigma(x)}p_{\epsilon}\left(\frac{y-\mu(x)}{\sigma(x)}\right).$
@@
$\hat{p}_{ls}(y|x)$ $=\frac{1}{\hat{\sigma}(x)}\hat{p}_{\epsilon}\left(\frac{y-\hat{\mu}(x)}{\hat{\sigma}(x)}\right).$
@@ 3.4 Implementation details
The kernel and local linear methods are implemented in the R package hdrcde by Hyndman et al. [14]. We use a Gaussian kernel and the bandwidth is selected based on the normal reference rule. The local kernel method utilizes the same package but with additional local moving windows. In our two case studies, we adopt exponentially growing window sizes to account for data sparsity in the high $p_{\perp}$ region. Specifically, we use windows of size $\exp(x/400)$ in the first case study (Sections 4 and 5) and $\exp(x/100)$ in the second case study (Section 6). For the location-scale model, the mean function $\mu$ and variance function $\sigma^{2}$ are estimated by smoothing splines using the core R function smooth.spline. The error distribution $\epsilon$ is estimated by kernel density estimation using the R function kde from the ks package [15]. The code for all implemented methods is available at https://github.com/richardzhs/response_matrix_estimation.
@@ 4.1 Simulation setup
$f\left(p_{\perp}\right)=LN_{0}\left(\frac{p_{\perp}}{\mathrm{GeV}}\right)^{-\alpha}\left(1-\frac{2}{\sqrt{s}}p_{\perp}\right)^{\beta}e^{-\gamma/p_{\perp}},$
@@
$\left(\frac{\sigma\left(p_{\perp}\right)}{p_{\perp}}\right)^{2}=\left(\frac{C_{1}}{\sqrt{p_{\perp}}}\right)^{2}+\left(\frac{C_{2}}{p_{\perp}}\right)^{2}+C_{3}^{2},$
@@
$g\left(p_{\perp}^{\prime}\right)=\int_{T}N\left(p_{\perp}^{\prime}\mid p_{\perp},\sigma\left(p_{\perp}\right)^{2}\right)f\left(p_{\perp}\right)\mathrm{d}p_{\perp},\quad p_{\perp}^{\prime}\in S.$
In this simulation, the particle-level and detector-level spaces are restricted to $T=S=[400\text{ GeV},1000\text{ GeV}]$. We partition both spaces using $m=n=40$ bins which corresponds to a response matrix of size $40\times 40$.
@@ 4.2 Results
Figure 2 shows the estimated response matrices for the different conditional density estimation methods described in Section 3. The sample size to estimate the response matrix is N=100, [^0] N=100{,}000. Pure eye inspection indicates that the histogram estimate becomes noticeably noisier toward the sparsely sampled right tail of the spectrum compared to the CDE methods.
@@ 5 Effect of the estimated response matrix on the unfolded spectrum
The sample size of the Monte Carlo sample for estimating the response matrix is the same as before with N=100, [^0] N=100{,}000. The sample size of the experimental data for constructing the histogram $\bm{y}$ is 1,000,000, mimicking a typical situation at the LHC where the available MC sample is smaller than the experimental sample. The performance of the estimators is compared using the bin-wise mean-squared error (MSE)
@@ 5.1 Numerical results for Tikhonov regularization
Figures 46 illustrate one realization of unfolded solutions with Tikhonov regularization using various regularization strengths $\delta$. Figures 79 show the corresponding (log-scale) mean squared error (MSE), squared bias and variance for M=1, [^0] M=1{,}000 realizations. In Figures 7 and 8, when $\delta=10^{-9}$ or $\delta=10^{-10}$, the solution using the true response matrix has the lowest MSE, followed by the location-scale method. In contrast, the histogram estimate exhibits the highest MSE. Notice that for the high $p_{\perp}$ region ($p_{\perp}\geq 600\text{ GeV}$), the MSE is largely driven by the variance (recall that $\text{MSE}=\text{bias}^{2}+\text{variance}$). The histogram method exhibits the largest variance, while the solution using the true matrix has the smallest variance. The location-scale method also has a relatively smaller variance compared to the other methods. This indicates that the variance of the Tikhonov-regularized solution is closely related to the quality of the estimated response matrix. Overall, we find that the relative performance of the unfolded point estimators aligns well with the relative performance of each response matrix estimation method in Section 4, supporting the hypothesis that a more accurately estimated response matrix leads to a better unfolded solution.
Figure 4: Tikhonov solution with $\delta=10^{-9}$ for the different response matrix estimators.
Figure 5: Tikhonov solution with $\delta=10^{-10}$ for the different response matrix estimators.
Figure 6: Tikhonov solution with $\delta=0$ (i.e., the unregularized least-squares solution) for the different response matrix estimators.
5.1.1 Noisy matrix acts as an implicit regularizer
An interesting observation is that when $\delta=0$ (Figures 6 and 9), i.e., without any regularization, the least-squares solution using the true response matrix performs surprisingly poorly. It has the largest MSE, bias and variance across most bins. Table 1 illustrates that the true response matrix has a notably larger condition number compared to the estimated response matrices. This explains the large variance of the solution with the true matrix. However, even with this extreme ill-conditioning, the least-squares solution should still, in principle, be unbiased. The fact that we nevertheless observe a fair amount of bias for it indicates that numerical artifacts when performing the ill-conditioned matrix inversion are large enough to create this nontrivial bias.
@@ 5.1.1 Noisy matrix acts as an implicit regularizer
Figure 7: Bin-wise mean squared error (left), squared bias (middle) and variance (right) for Tikhonov regularization with $\delta=10^{-9}$.
Figure 8: Bin-wise mean squared error (left), squared bias (middle) and variance (right) for Tikhonov regularization with $\delta=10^{-10}$.
Figure 9: Bin-wise mean squared error (left), squared bias (middle) and variance (right) for Tikhonov regularization with $\delta=0$ (i.e., the unregularized least-squares solution).
@@ 5.2 Numerical results for D’Agostini iteration
Figures 1012 illustrate one D’Agostini solution with various numbers of iterations for the different response matrix estimators. Figures 1315 show the corresponding (log-scale) MSE, squared bias and variance for M=1, [^0] M=1{,}000 realizations. As in the case of Tikhonov regularization, the MSE of the D’Agostini solutions is also mostly driven by the variance. The relative magnitude of the variance aligns with the performance of each response matrix estimation method, with a better estimated response matrix generally leading to lower variance. But the difference in the variances is not as significant as in the Tikhonov case when the number of iterations is modest (Figures 13 and 14). This indicates that the statistical uncertainty from the estimated response matrix is overshadowed by the strong regularization in this case. However, as the number of iterations increases (Figure 15), we see that the difference in the variances becomes more pronounced.
Figure 10: D’Agostini solution with 5 iterations for the different response matrix estimators.
Figure 11: D’Agostini solution with 30 iterations for the different response matrix estimators.
Finally, in both the Tikhonov and D’Agostini solutions, the bias of the global kernel method is significantly higher than for the other methods in the low $p_{\perp}$ region. This is because the estimated response matrix has a large bias in this region, as observed in Section 4. As we discussed above, there does not seem to exist a single pair of global bandwidths that can produce a good estimate for all $p_{\perp}$. In this case, the chosen bandwidths oversmooth in the low $p_{\perp}$ region, resulting in a high bias for the estimated matrix which is then also reflected in the unfolded point estimate. Similarly, for the local linear method, the variance is larger than for the other methods (except for the naive histogram method) in the high $p_{\perp}$ region. This is due to the fact that the chosen bandwidths undersmooth in the high $p_{\perp}$ region. On the other hand, by using local bandwidths, the local kernel method provides a “best-of-both-worlds” approach by having, in most cases, bias comparable to the local linear method and variance comparable to the global kernel method.
Figure 12: D’Agostini solution with 5,000 iterations for the different response matrix estimators.
Figure 13: Bin-wise mean squared error (left), squared bias (middle) and variance (right) for D’Agostini unfolding with 5 iterations.
Figure 14: Bin-wise mean squared error (left), squared bias (middle) and variance (right) for D’Agostini unfolding with 30 iterations.
Figure 15: Bin-wise mean squared error (left), squared bias (middle) and variance (right) for D’Agostini unfolding with 5,000 iterations.

### v2

@@ Abstract
The unfolding problem in particle physics is to make inferences about the true particle spectrum based on smeared observations from a detector. This is an ill-posed inverse problem, where small changes in the smeared distribution can lead to large fluctuations in the unfolded distribution. The forward operator is the response matrix which models the detector response. In practice, the forward operator is rarely known analytically and is instead estimated using Monte Carlo simulation. This raises the question of how to best estimate the response matrix and what impact this estimation has on the unfolded solutions. In most analyses at the LHC, response matrix estimation is done by binning the true and smeared events and counting the propagation of events between the bins. However, this approach can result in a noisy estimate of the response matrix, especially with a small Monte Carlo sample size. Unexpectedly, we also find that the noise in the estimated response matrix can inadvertently regularize the problem. As an alternative, we propose to estimate the response matrix through the use of conditional density estimation of the response kernel in the unbinned setting followed by binning this estimator. Using simulation studies, we investigate the performance of the two approaches.
@@
Analysis and statistical methods, Data analysis, Large detector systems for particle and astroparticle physics
journal: Nuclear Instruments and Methods in Physics Research A
@@ 1 Introduction
The unfolding problem arises in the statistical analysis of data at the Large Hadron Collider (LHC) at CERN. In LHC experiments, it is often of interest to learn the distribution (or differential cross section) $f$ of some physical quantity measured by the detectors, e.g., the energy, mass or production angle of the particles produced in the proton-proton collisions. Due to the finite resolution of the detectors, however, the observed data follow a smeared distribution $g$, rather than $f$ itself. This detector smearing maps the true distribution $f$ to the measured distribution $g$, as illustrated in Figure 1. Unfolding refers to the process of inferring $f$ from $g$ (for details, see, e.g., [7]).
@@
In practice, the response matrix $\bm{K}$ is usually not known analytically and needs to be estimated using detector simulations. This raises the question of how to best estimate the response matrix and what impact this estimation has on the unfolded solutions. The main contribution of this work is to introduce a new way to estimate the response matrix by first considering response kernel estimation in the unbinned setting. Additionally, we show that there are nontrivial, and sometimes unexpected, interactions between the response matrix estimators and the quality of the unfolded solutions.
@@ 2.1 Discretization
$p_{X}(x)=\frac{f(x)}{\lambda(T)},\;\;p_{Y}(y)=\frac{g(y)}{\mu(S)},$ (2)
@@ 2.2 Unfolded point estimation
$\bm{L}^{-1}\mathbf{y}$ $=\bm{L}^{-1}\bm{K}\bm{\lambda}+\bm{\eta}.$ (8)
Since $\bm{\eta}\sim N(\bm{0},\bm{L}^{-1}\bm{\Sigma}(\bm{L}^{-1})^{\top})$ and $\bm{L}^{-1}\bm{\Sigma}(\bm{L}^{-1})^{\top}=\bm{I}$, it follows that $\bm{\eta}\sim N(\bm{0},\bm{I})$. Therefore, since $\bm{\mu}$ is accurately estimable using $\mathbf{y}$ and the Cholesky decomposition always exists for the covariance matrix, we may replace $\bm{K}\hookrightarrow\bm{L}^{-1}\bm{K}$ and assume without loss of generality that $\bm{\epsilon}\sim N(\bm{0},\bm{I})$. Under this assumption and assuming that $\bm{K}$ has full column rank, the least squares estimator of $\bm{\lambda}$ is
$\hat{\bm{\lambda}}_{ls}=(\bm{K}^{\top}\bm{K})^{-1}\bm{K}^{\top}\mathbf{y}.$ (9)
@@
$\mathbb{E}[\hat{\bm{\lambda}}_{ls}]=(\bm{K}^{\top}\bm{K})^{-1}\bm{K}^{\top}\bm{K}\bm{\lambda}=\bm{\lambda}.$ (10)
@@
$\hat{\bm{\lambda}}_{tik}$ $=\text{argmin}_{\bm{\lambda}}\left\{\|\mathbf{y}-\bm{K}\bm{\lambda}\|_{2}^{2}+\delta\|\bm{\lambda}\|^{2}_{2}\right\}$ (11)
$=(\bm{K}^{\top}\bm{K}+\delta\bm{I})^{-1}\bm{K}^{\top}\mathbf{y},$
@@ 3.2.1 Kernel regression method
Estimating the conditional density can be reduced to a non-parametric regression problem. Let $K(\cdot)$ be a symmetric non-negative density function. Denote $f(y|x):=\mathbb{E}[K_{h}(Y-y)|X=x]$, where $K_{h}(\cdot)=K(\cdot/h)/h$ and $h>0$ is the bandwidth parameter that controls the degree of smoothing (see, e.g., [30, 28]). Then, by Taylor expansion,
$f(y_{0}|x)$ $=\int K_{h}(y^{\prime}-y_{0})p_{Y|X}(y^{\prime}|x)dy^{\prime}$ (14)
@@ 3.2.2 Local linear method
This yields the local linear estimate $\hat{p}_{ll}(y|x)=\hat{f}_{h_{1},h_{2}}(y|x)$. Note that if we set $b=0$ in the objective function (18), then the solution to the minimization problem recovers the kernel conditional density estimate $\hat{p}_{ks}(y|x)$. Compared to the kernel method, one attractive feature of the local linear method is its design adaptive property which states that asymptotically the bias of $\hat{p}_{ll}(y|x)$ does not depend on the marginal distribution of $X$ [16]. This property is particularly important because the distribution of $X$ or its parametric form is not known a priori. Moreover, in practice, the marginal distribution may fall steeply over several orders of magnitude.
@@ 3.2.3 Bandwidth selection
There are many existing methods for picking $h_{1}$ and $h_{2}$. One of the simplest and computationally most efficient ways is to find the optimal bandwidths by assuming the normality of both the conditional and marginal distributions [11, 20]. Although it relies on the normality assumption, this can yield reasonable results even in cases where normality does not hold [28]. We also find this works reasonably well in our examples.
@@ 3.2.5 Location-scale model
$p_{Y|X}(y|x)=\frac{1}{\sigma(x)}p_{\epsilon}\left(\frac{y-\mu(x)}{\sigma(x)}\right).$ (24)
@@
$\hat{p}_{ls}(y|x)$ $=\frac{1}{\hat{\sigma}(x)}\hat{p}_{\epsilon}\left(\frac{y-\hat{\mu}(x)}{\hat{\sigma}(x)}\right).$ (25)
@@ 3.4 Implementation details
The kernel and local linear methods are implemented in the R package hdrcde by Hyndman et al. [19]. We use a Gaussian kernel and the bandwidth is selected based on the normal reference rule. The local kernel method utilizes the same package but with additional local moving windows. In our two case studies, we adopt exponentially growing window sizes to account for data sparsity in the high $p_{\perp}$ region. Specifically, we use windows of size $\exp(x/400)$ in the first case study (Sections 4 and 5) and $\exp(x/100)$ in the second case study (Section 6). The window sizes differ slightly to reflect the different kinematic ranges of the two studies: 400–1000 GeV for the first case study and 20–400 GeV for the second. For the location-scale model, the mean function $\mu$ and variance function $\sigma^{2}$ are estimated by smoothing splines using the core R function smooth.spline. The error distribution $\epsilon$ is estimated by kernel density estimation using the R function kde from the ks package [14]. The code for all implemented methods is available at https://github.com/richardzhs/response_matrix_estimation.
@@ 4.1 Simulation setup
$f\left(p_{\perp}\right)=LN_{0}\left(\frac{p_{\perp}}{\mathrm{GeV}}\right)^{-\alpha}\left(1-\frac{2}{\sqrt{s}}p_{\perp}\right)^{\beta}e^{-\gamma/p_{\perp}},$ (27)
@@
$\left(\frac{\sigma\left(p_{\perp}\right)}{p_{\perp}}\right)^{2}=\left(\frac{C_{1}}{\sqrt{p_{\perp}}}\right)^{2}+\left(\frac{C_{2}}{p_{\perp}}\right)^{2}+C_{3}^{2},$ (28)
@@
$g\left(p_{\perp}^{\prime}\right)=\int_{T}N\left(p_{\perp}^{\prime}\mid p_{\perp},\sigma\left(p_{\perp}\right)^{2}\right)f\left(p_{\perp}\right)\mathrm{d}p_{\perp},\quad p_{\perp}^{\prime}\in S.$ (29)
In this simulation, the particle-level and detector-level spaces are restricted to $T=S=[400\text{ GeV},1000\text{ GeV}]$. We partition both spaces using $m=n=40$ bins which corresponds to a response matrix of size $40\times 40$. The 7 TeV case study has been used in several previous methodological studies [22, 23], allowing for a direct comparison with those works. However, to reflect more current experimental conditions, we also consider a 13 TeV case study in Section 6.
@@ 4.2 Results
Figure 2 shows the estimated response matrices for the following methods: (1) histogram: direct estimation of the response matrix via events binning (Section 3.1); (2) global kernel: kernel conditional density estimation with global bandwidths (Section 3.2.1); (3) local linear: local linear conditional density estimation with global bandwidths (Section 3.2.2); (4) local kernel: kernel conditional density estimation with locally adaptive bandwidths (Section 3.2.4); (5) location-scale: estimation of mean and variance functions assuming a location-scale model (Section 3.2.5); and (6) true response matrix (for reference). The sample size to estimate the response matrix is N=100, [^0] N=100{,}000. Pure eye inspection indicates that the histogram estimate becomes noticeably noisier toward the sparsely sampled right tail of the spectrum compared to the CDE methods.
@@ 5 Effect of the estimated response matrix on the unfolded spectrum
The sample size of the Monte Carlo sample for estimating the response matrix is the same as before with N=100, [^0] N=100{,}000. The sample size of the experimental data for constructing the histogram $\mathbf{y}$ is 1,000,000, mimicking those situations at the LHC where the available MC sample is smaller than the experimental sample. The performance of the estimators is compared using the bin-wise mean-squared error (MSE)
@@ 5.1 Numerical results for Tikhonov regularization
(a) Tikhonov solution with $\delta=10^{-9}$.
(b) Tikhonov solution with $\delta=10^{-10}$.
Figure 4: Tikhonov solutions with regularization parameters (a) $\delta=10^{-9}$ and (b) $\delta=10^{-10}$.
Figures 45 illustrate one realization of unfolded solutions with Tikhonov regularization using various regularization strengths $\delta$. Figure 6 shows the corresponding (log-scale) mean squared error (MSE), squared bias and variance for M=1, [^0] M=1{,}000 realizations. Rather than selecting $\delta$ according to an optimality criterion such as MSE or goodness-of-fit statistic, we intentionally consider a broad range of values, including $\delta=0$ (i.e. no regularization), in order to assess how the unfolded solution behaves across different regularization regimes. In Figures 6(a) and 6(b), when $\delta=10^{-9}$ or $\delta=10^{-10}$, the solution using the true response matrix generally has the lowest MSE, followed by the location-scale method. In contrast, the histogram estimate exhibits the highest MSE. Notice that for the high $p_{\perp}$ region ($p_{\perp}\geq 600\text{ GeV}$), the MSE is largely driven by the variance (recall that $\text{MSE}=\text{bias}^{2}+\text{variance}$). In this regime, the histogram method shows the largest variance, whereas the solution based on the true response matrix has the smallest variance, with the location–scale method also exhibiting relatively low variance compared to the other approaches. This behavior indicates that the variance of the Tikhonov-regularized solution is closely related to the quality of the estimated response matrix. Overall, we find that the relative performance of the unfolded point estimators aligns well with the relative performance of each response matrix estimation method in Section 4, supporting the hypothesis that a more accurately estimated response matrix leads to a better unfolded solution.
@@
Figure 5: Unregularized least-squares solution ($\delta=0$).
5.1.1 Noisy matrix acts as an implicit regularizer
An interesting observation is that when $\delta=0$ (Figures 5 and 6(c)), i.e., without any regularization, the least-squares solution using the true response matrix performs surprisingly poorly. It has the largest MSE, bias and variance across most bins. Table 1 illustrates that the true response matrix has a notably larger condition number compared to the estimated response matrices. This explains the large variance of the solution with the true matrix. However, even with this extreme ill-conditioning, the least-squares solution should still, in principle, be unbiased. The non-negligible bias we nevertheless observe may be indicative of numerical artifacts arising from the ill-conditioned matrix inversion.
(a) $\delta=10^{-9}$.
(b) $\delta=10^{-10}$.
(c) $\delta=0$ (unregularized solution).
Figure 6: Bin-wise mean squared error (left), squared bias (middle), and variance (right) for Tikhonov regularization with different $\delta$ values. All three panels for each $\delta$ share the same y-axis for direct comparison.
@@ 5.1.1 Noisy matrix acts as an implicit regularizer
(a) D’Agostini unfolding solution with 55 iterations.
(b) D’Agostini unfolding solution with $30$ iterations.
Figure 7: D’Agostini solutions with (a) 55 iterations and (b) $30$ iterations.
Figure 8: D’Agostini solution with 5,0005{,}000 iterations.
@@ 5.2 Numerical results for D’Agostini iteration
Figures 78 illustrate one D’Agostini solution with various numbers of iterations for the different response matrix estimators. Figure 9 shows the corresponding (log-scale) MSE, squared bias and variance for M=1, [^0] M=1{,}000 realizations. As in the case of Tikhonov regularization, the MSE of the D’Agostini solutions is also mostly driven by the variance. The relative magnitude of the variance aligns with the performance of each response matrix estimation method, with a better estimated response matrix generally leading to lower variance. But the difference in the variances is not as significant as in the Tikhonov case when the number of iterations is modest (Figures 9(a) and 9(b)). This indicates that the statistical uncertainty from the estimated response matrix is overshadowed by the strong regularization in this case. However, as the number of iterations increases (Figure 9(c)), we see that the difference in the variances becomes more pronounced.
In both the Tikhonov and D’Agostini solutions, the bias of the global kernel method is significantly higher than for the other methods in the low $p_{\perp}$ region. This is because the estimated response matrix has a large bias in this region, as observed in Section 4. As we discussed above, there does not seem to exist a single pair of global bandwidths that can produce a good estimate for all $p_{\perp}$. In this case, the chosen bandwidths oversmooth in the low $p_{\perp}$ region, resulting in a high bias for the estimated matrix which is then also reflected in the unfolded point estimate. Similarly, for the local linear method, the variance is larger than for the other methods (except for the naive histogram method) in the high $p_{\perp}$ region. This is due to the fact that the chosen bandwidths undersmooth in the high $p_{\perp}$ region. On the other hand, by using local bandwidths, the local kernel method provides a “best-of-both-worlds” approach by having, in most cases, bias comparable to the local linear method and variance comparable to the global kernel method.
(a) $niter=5$.
(b) $niter=30$.
(c) $niter=5000$.
Figure 9: Bin-wise mean squared error (left), squared bias (middle), and variance (right) for D’Agostini solutions with different iteration counts. All three panels for each niterniter share the same y-axis for direct comparison.